OpenStudy (anonymous):

evaluate given a=-3, b=4, c=-5 a^2-abs bc abs

OpenStudy (anonymous):

i got -29 how did you get the 60

OpenStudy (anonymous):

why are we sub (-3) with the 4 and 5

OpenStudy (anonymous):

it say a^2-\[\left| bc \right|\] a=-3 b=4 c=-5

OpenStudy (anonymous):

I am not sure that is why I'm confused I guess.

OpenStudy (anonymous):

I got this wrong on a test and was trying to figure out what I did wrong is all

OpenStudy (anonymous):

it states a=-3 b=4 and c=-5 \[a ^{2}-\left| bc \right|\]

OpenStudy (anonymous):

evaluate

OpenStudy (anonymous):

you just drop the abs value signs and add

OpenStudy (anonymous):

sorry subtract

OpenStudy (anonymous):

i don't know I'm confused

OpenStudy (anonymous):

Thanks